# Procountable groups are not classifiable by countable structures

### v2

@@ Abstract.
We prove that topological isomorphism on $2$-nilpotent procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ that expresses that two sequences of reals have bounded difference is Borel reducible to it. This marks substantial progress on an open problem of [16]: to determine the exact complexity of the isomorphism relation among all non-archimedean Polish groups.
S. Gao acknowledges the partial support of their research by the Fundamental Research Funds for the Central Universities and by the National Natural Science Foundation of China (NSFC) grant 12271263. A. Nies acknowledges support by the Marsden Fund of New Zealand, UOA-1931. Both S. Gao and A. Nies attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this research was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center. G. Paolini was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA and by INdAM Project 2024 (consolidator grant) “Groups, Crystals and Classifications”.
@@ 1. Introduction
It is often useful to have concrete presentations of abstract mathematical objects of a certain type. In this case one is interested in the complexity of whether two presentations present the same abstract object. For instance, finite graphs can be concretely presented by adjacency matrices (written as binary strings). The graph isomorphism problem asks whether two such matrices present the same graph. This problem is in NP, and its exact complexity has been the object of much recent study [1, 31]. The triviality problem for groups is to determine whether a presentation given by finitely many generators and relators presents the trivial group. This is undecidable [30], so the $\Sigma^{0}_{1}$ relation of isomorphism between finite presentations of groups is undecidable. In fact, every $\Sigma^{0}_{1}$ equivalence relation is computably reducible to it by [23] (also see [26, Cor. 2.6]).
For many types of mathematical objects, a countably infinite amount of information is required to present a member. In this case, one can usually view the presentations as points in a so-called standard Borel space (such as the unit interval). Borel reducibility, introduced in [7], is the standard tool for comparing the complexity of equivalence relations on standard Borel spaces. For instance, countable graphs can be concretely presented as symmetric, irreflexive relations on $\mathbb{N}$, and their isomorphism is a well-known benchmark equivalence relation [8] denoted by GI. The isomorphism relation for a class of concretely presented objects is reducible to GI if and only there is a Borel assignment of countable structures as complete invariants to presentations of the objects in question.
A programme of Kechris et al. A Polish group is a topological group with a Polish topology, namely, one that is separable and completely metrizable. A topological group is called non-archimedean if it has a neighbourhood basis of the identity consisting of open subgroups [2]. Such groups are totally disconnected. Let $\mathrm{Sym}(\mathbb{N})$ denote the group of all permutations of $\mathbb{N}$ equipped with the pointwise convergence topology. Then $\mathrm{Sym}(\mathbb{N})$ is a non-archimedean Polish group. Up to topological isomorphism, the non-archimedean Polish groups are precisely the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ [2]. It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space (see [16]). So we can take the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ as the concrete presentations of non-archimedean Polish groups.
It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space; see for instance Kechris et al. [16]. In that article they initiated the following programme:
@@
There has been some progress on this program. We summarize the results known to us. One class considered in [16] is the class of compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bi-reducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel. One can also consider the subclass of abelian profinite groups. Due to Pontryagin duality, the isomorphism relation between abelian profinite groups is Borel bi-reducible with the isomorphism relation between countable torsion abelian groups. The latter relation is classified by Ulm invariants; it is also a complete analytic set and hence non-Borel (see [7, (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$. The isomorphism relation of oligomorphic groups was shown to be essentially countable in [24]; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see [25, 27]; in particular in [27] it was shown that under the assumption of weak elimination of imaginaries the relation is smooth. The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in [4], where it is shown that the setwise equality of infinite sequences of reals is Borel reducible to it. Work in progress of Ruiwen Li, in conjunction with [4], would show that in fact GI is reducible to it.
We note that the programme of [16] asked to first determine whether a given class of interest $\mathcal{C}$ of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ is indeed Borel. While this is often straightforward from the definition, in some cases it takes work to establish. Borelness is unknown for the class of topologically finitely generated groups, and known to fail for some classes of interest. For instance, the class of all CLI groups (those admitting a compatible complete left-invariant metric) was shown to be properly co-analytic by Malicki [19]. Paolini and Shelah [29] proved that four properties of closed abelian subgroups of $\mathrm{Sym}(\mathbb{N})$ are complete co-analytic and therefore non-Borel: separability (in the sense of abelian group theory), torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
Countable groups are up to isomorphism the discrete closed subgroups $G$ of $\mathrm{Sym}(\mathbb{N})$, and being discrete is Borel because it means that the identity permutation is isolated in $G$. If a class of discrete subgroups of $\mathrm{Sym}(\mathbb{N})$ is Borel, its isomorphism relation is Borel reducible to GI. The question is whether conversely, GI reduces to the isomorphism relation on the class. Mekler [22] (as noted in [7, 2.3]) established this for the countable groups that are nilpotent of class $2$, and of exponent a fixed odd prime. His method of coding graphs into such groups will be instrumental for our main result, Theorem 1.2. Paolini and Shelah [28] showed it for countable torsion-free abelian groups. Most recently, Gao and Li [10] did so for countable omnigenous locally finite groups (an uncountable class of groups, each generalising Hall’s universal locally finite group).
The main result. Our work marks considerable progress on an open problem of Kechris et al. [16] pertaining to their programme. They asked at the end of their introduction for the complexity of isomorphism on the class of all non-archimedean Polish groups, observing that this relation is analytic, and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of 2-nilpotent procountable groups of exponent a fixed odd prime $p$ already has a complexity much higher than GI.
@@
It is easily seen that a group of the form $\prod_{n}G_{n}$, and hence each procountable group, is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. For background and further equivalent characterisations, see e.g., [20]. We will verify in 3.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of the neutral element consisting of normal open subgroups (see, e.g., [12, Theorem 3.3] or [20, Lemma 2]). The procountable groups satisfying some law (such as being 2-nilpotent) then also form a Borel set: from a closed subgroup $G$ of $\mathrm{Sym}(\mathbb{N})$, using [14, Th. 12.13] one can in a Borel way determine a countable dense subgroup $D$, and it satisfies to require that $D$ satisfies the law.
Since each countable discrete group is procountable, by the aforementioned result of Mekler [22], GI is reducible to the isomorphism relation on 2-nilpotent procountable groups of exponent $p$. It follows that the topological isomorphism relation of procountable groups is a complete analytic set and hence non-Borel.
Recall that $E_{1}$ is the Borel equivalence relation of eventual equality of sequences of real numbers. By a celebrated theorem of Kechris and Louveau [15], $E_{1}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. It follows that any equivalence relation to which $E_{1}$ is Borel reducible has the same property. Recall that $\ell_{\infty}$ is the Borel equivalence relation on $\mathbb{R}^{\mathbb{N}}$ given by
@@
By a result of Rosendal (see [8, Theorem 8.4.2]), $E_{1}$ is Borel reducible to $\ell_{\infty}$, and thus $\ell_{\infty}$ is not Borel reducible to the orbit equivalence relation of any Borel action of a Polish group. The following is the main result of this paper.
@@
The proof makes use of the notion of uniform homeomorphism.
@@
Open questions.
We close the introduction with some open questions. The first two will be discussed again in see Section 2.4. An analytic equivalence relation on a Borel space is called universal analytic if each analytic equivalence relation is Borel reducible to it.
Question 1.4.
@@
We note that our original construction for the second step used as $G_{n}$ the free Coxeter group generated by $V_{n}$, namely, the group freely generated by involutions $g_{v}$ for $v\in V_{n}$; see the arxiv version 2512.12256.v1. That approach does not yield nilpotent-2 groups. However, it is somewhat simpler to verify, because the only elements of order 2 in $G_{n}$ are the conjugates of the generators, so that the elements of $V_{n}$ up to conjugacy can be recovered from $G_{n}$. It is possible that this method could lead to an affirmative answer Question 1.4.
It is known that the relation of uniform homeomorphism on Polish metric spaces is universal analytic [5].
Question 1.5.
Is the equivalence relation of uniform homeomorphism on ultrametric Polish spaces universal analytic?
By the second step above, an affirmative answer to the second question would also answer the first in the affirmative.
@@
Notice that it follows from [28] that GI is a lower bound for the complexity of this classification problem.
@@ 2. Borel reduction of $\ell_{\infty}$ to uniform homeomorphism
We use the main result of [3], where it was shown that GI is Borel bireducible with the homeomorphism relation between 0-dimensional compact metric spaces. Let $\Phi$ denote a Borel witness to this reduction. Then for any graphs $\Gamma,\Gamma^{\prime}$ with vertex set $\mathbb{N}$, $\Gamma$ and $\Gamma^{\prime}$ are isomorphic if and only if $\Phi(\Gamma)$ and $\Phi(\Gamma^{\prime})$ are homeomorphic. Note that a Polish space is 0-dimensional if and only if it admits a compatible ultrametric. So there exists a compatible ultrametric $d_{\Gamma}$ for $\Phi(\Gamma)$. By the standard tools from descriptive set theory, we may find a Borel map giving the correspondence $\Phi(\Gamma)\mapsto d_{\Gamma}$.
@@ 2.1. Some facts on uniform homeomorphisms
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$. Then $\Phi$ is uniformly continuous if and only if for all $n\in\mathbb{N}$ there is $m\in\mathbb{N}$ such that for any $x,y\in[T]$, if $x\!\!\upharpoonright\!m=y\!\!\upharpoonright\!m$, then $\Phi(x)\!\!\upharpoonright\!n=\Phi(y)\!\!\upharpoonright\!n$.
@@
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$ be a uniform homeomorphism (see 1.3). If both $[T],[S]$ are uncountable, then $\Omega_{\Phi}$ and $\Omega_{\Phi^{-1}}$ are unbounded monotone functions from $\mathbb{N}$ to $\mathbb{N}$.
@@ 2.2. Definition of the reduction
The trees $\mathsf{S}_{1},\mathsf{S}_{2},\dots,\mathsf{S}_{k},\dots$ will be abstractly and inductively defined. Their exact realizations as subtrees of $\mathbb{N}^{<\mathbb{N}}$ will not be important. To begin, let $\mathsf{S}_{1}$ be a single branch. Thus the Cantor–Bendixson rank of $\mathsf{S}_{1}]$ is $1$. Let $\mathsf{S}_{2}$ be a pruned tree so that the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $\mathsf{S}_{1}]$. So the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $2$. In general, let $\mathsf{S}_{k+1}$ be a pruned tree whose Cantor–Bendixson derivative is $\mathsf{S}_{k}]$. Then the Cantor–Bendixson rank of $\mathsf{S}_{k+1}$ is $k+1$. This finishes the definitions of $\mathsf{S}_{k}$ for $1\leqslant k<\mathbb{N}$.
@@
for any clopen set $U$ with $z\in U$, the $k$-th Cantor–Bendixson derivative of $U$ is the perfect kernel of $U$, and for any $i<k$, the $i$-th Cantor–Bendixson derivative of $U$ contains a perfect subset but is not itself perfect.
@@
Finally we are ready to define $\mathsf{T}_{x}$ given an $x\in\mathbb{N}^{\mathbb{N}}$. $\mathsf{T}_{x}$ is obtained from $\mathsf{S}_{\omega}$ by extending the node $\mathsf{t}_{k,2m+1}$ by a copy of $\mathsf{R}_{x(k)}$ for all $k,m\in\mathbb{N}$. This finishes the definition of $\mathsf{T}_{x}$.

### v3

@@ Abstract.
We prove that topological isomorphism on $2$-nilpotent procountable groups is not classifiable by countable structures, in the sense of descriptive set theory. In fact, the equivalence relation $\ell_{\infty}$ expressing that two sequences of reals have a bounded difference is Borel reducible to it. This marks substantial progress on an open problem of Kechris, Nies and Tent (2018): to determine the exact complexity of the isomorphism relation among all non-archimedean Polish groups.
S. Gao acknowledges the partial support of their research by the Fundamental Research Funds for the Central Universities and by the National Natural Science Foundation of China (NSFC) grant 12271263. A. Nies acknowledges support by the Marsden Fund of New Zealand, UOA-346. Both S. Gao and A. Nies attended the Tianyuan Workshop on Computability Theory and Descriptive Set Theory in June 2025 when a part of this research was conducted, and we acknowledge the support by the Tianyuan Mathematics Research Center. G. Paolini was supported by project PRIN 2022 “Models, sets and classifications”, prot. 2022TECZJA and by INdAM Project 2024 (consolidator grant) “Groups, Crystals and Classifications”.
@@ 1. Introduction
It is often useful to have concrete presentations of abstract mathematical objects of a certain type. In this case one is interested in the complexity of whether two presentations present the same abstract object. For instance, finite graphs can be concretely presented by adjacency matrices (written as binary strings). The graph isomorphism problem asks whether two such matrices present the same graph. This problem is in NP, and its exact complexity has been the object of much recent study Babai2016 ; Wiebking2019 . For many types of mathematical objects, a countably infinite amount of information is required to present a member. In this case, one can usually view the presentations as points in a so-called standard Borel space (such as the unit interval). Borel reducibility, introduced in FriedmanStanley1989 , is the standard tool for comparing the complexity of equivalence relations on standard Borel spaces. For instance, countable graphs can be concretely presented as symmetric, irreflexive relations on $\mathbb{N}$, and their isomorphism is a well-known benchmark equivalence relation GaoBook denoted by GI. The isomorphism relation for a class of concretely presented objects is reducible to GI if and only there is a Borel assignment of countable structures as complete invariants to presentations of the objects in question.
The main result. A Polish group is a topological group with a Polish topology, namely, one that is separable and completely metrizable. A topological group is called non-archimedean if it has a neighbourhood basis of the identity consisting of open subgroups BeckerKechris1996 . Such groups are totally disconnected. Let $\mathrm{Sym}(\mathbb{N})$ denote the group of all permutations of $\mathbb{N}$ equipped with the pointwise convergence topology. Then $\mathrm{Sym}(\mathbb{N})$ is a non-archimedean Polish group. Up to topological isomorphism, the non-archimedean Polish groups are precisely the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ BeckerKechris1996 . It is well known that the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a standard Borel space; see for instance Kechris et al. KNT . So the closed subgroups of $\mathrm{Sym}(\mathbb{N})$ are concrete presentation of non-archimedean Polish groups.
In the same article KNT , the authors initiated the following programme:
@@
Our work marks considerable progress on an open problem of Kechris et al. KNT pertaining to their programme. They asked at the end of their introduction for the complexity of isomorphism on the class of all non-archimedean Polish groups, observing that this relation is analytic, and GI is Borel reducible to it. By our main result, Theorem 1.2 below, the isomorphism relation on the Borel class of 2-nilpotent procountable groups of exponent a fixed odd prime $p$ already has a complexity much higher than GI.
@@
Procountable groups are of interest because they generalise both profinite and discrete groups, and on the other hand they are precisely the non-archimedean groups that have small invariant neighbourhoods (SIN). (See Pestov:99 for a discussion of this important class of topological groups.) It is easily seen that a group of the form $\prod_{n}G_{n}$, and hence each procountable group, is isomorphic to a closed subgroup of $\mathrm{Sym}(\mathbb{N})$. For background and further equivalent characterisations, see e.g., Malicki2016 . We will verify in 3.3 that the procountable closed subgroups of $\mathrm{Sym}(\mathbb{N})$ form a Borel set, using that the procountable groups are the non-archimedean groups with a neighbourhood basis of the neutral element consisting of normal open subgroups (see, e.g., [GX, , Theorem 3.3] or [Malicki2016, , Lemma 2]). The procountable groups satisfying some law (such as being 2-nilpotent) also form a Borel set: from a closed subgroup $G$ of $\mathrm{Sym}(\mathbb{N})$, using [Kechris, , Th. 12.13] one can in a Borel way determine a countable dense subgroup $D$, and it satisfies to require that $D$ satisfies the law.
Since each countable discrete group is procountable, by the aforementioned result of Mekler Mekler1985 , GI is reducible to the isomorphism relation on 2-nilpotent procountable groups of exponent $p$. It follows that the topological isomorphism relation of procountable groups is a complete analytic set and hence non-Borel.
Recall that $E_{1}$ is the Borel equivalence relation of eventual equality of sequences of real numbers. By a celebrated theorem of Kechris and Louveau KL , $E_{1}$ is not Borel reducible to any orbit equivalence relation induced by a Borel action of a Polish group. It follows that any equivalence relation to which $E_{1}$ is Borel reducible has the same property. Recall that $\ell_{\infty}$ is the Borel equivalence relation on $\mathbb{R}^{\mathbb{N}}$ given by
@@
By a result of Rosendal (see [GaoBook, , Theorem 8.4.2]), $E_{1}$ is Borel reducible to $\ell_{\infty}$, and thus $\ell_{\infty}$ is not Borel reducible to the orbit equivalence relation of any Borel action of a Polish group. The following is the main result of this paper.
@@
Kechris et al. [KNT, , Theorem 3.1] provided a criterion when the topological isomorphism relation on a class $\mathcal{C}$ is Borel reducible to GI; intuitively, it says that one can canonically and in a Borel way assign to $G\in\mathcal{C}$ a countable neighbourhood basis of the identity consisting of open subgroups. By our main theorem, the hypothesis of this criterion is violated in our setting:
Corollary 1.3.
For an odd prime $p$, let $\mathcal{G}_{2,p}$ be the Borel class of all $2$-nilpotent procountable groups of exponent $p$. There is no assignment $G\mapsto\mathcal{N}_{G}$ such that for each $G\in\mathcal{G}_{2,p}$, $\mathcal{N}_{G}$ is a countable set of open subgroups of $G$ that form a neighborhood basis of $1_{G}$, and this assignment is Borel and isomorphism invariant in the sense of [KNT, , Theorem 3.1].
The proof of our main theorem makes use of the notion of uniform homeomorphism.
@@
Some background on the program of KNT . For context we summarize other results known to us that pertain to the program. One class considered in KNT is the class of compact non-archimedean Polish groups, or equivalently, the profinite groups. It was shown that its isomorphism relation is Borel bi-reducible with GI. This implies that this isomorphism relation is a complete analytic set, and hence non-Borel. One can also consider the subclass of abelian profinite groups. Due to Pontryagin duality, the isomorphism relation between abelian profinite groups is Borel bi-reducible with the isomorphism relation between countable torsion abelian groups. The latter relation is classified by Ulm invariants; it is also a complete analytic set and hence non-Borel (see [FriedmanStanley1989, , (2.1)]). In terms of Borel reducibility, it is strictly above the smooth equivalence relation and strictly below GI. In particular, it is not above the equivalence relation $E_{0}$. The isomorphism relation of oligomorphic groups was shown to be essentially countable in NiesSchlichtTent2022 ; it is at present an open question whether this relation is smooth (i.e., real numbers can be assigned as invariants); for partial results on this problem, see NP_oligo ; reconstruction2 ; in particular in reconstruction2 it was shown that under the assumption of weak elimination of imaginaries the relation is smooth. The isomorphism relation of extremely amenable non-archimedean Polish groups was studied in EGL , where it is shown that the setwise equality of infinite sequences of reals is Borel reducible to it. Work in progress of Ruiwen Li, in conjunction with EGL , would show that in fact GI is reducible to it.
We note that the programme of KNT asked to first determine whether a given class of interest $\mathcal{C}$ of closed subgroups of $\mathrm{Sym}(\mathbb{N})$ is indeed Borel. While this is often straightforward from the definition, in some cases it takes work to establish. Borelness is unknown for the class of topologically finitely generated groups, and known to fail for some classes of interest. For instance, the class of all CLI groups (those admitting a compatible complete left-invariant metric) was shown to be properly co-analytic by Malicki Ma . Paolini and Shelah PS_polish proved that four properties of closed abelian subgroups of $\mathrm{Sym}(\mathbb{N})$ are complete co-analytic and therefore non-Borel: separability (in the sense of abelian group theory), torsionlessness, $\aleph_{1}$-freeness and $\mathbb{Z}$-homogeneity.
Countable groups are up to isomorphism the discrete closed subgroups $G$ of $\mathrm{Sym}(\mathbb{N})$, and being discrete is Borel because it means that the identity permutation is isolated in $G$. If a class of discrete subgroups of $\mathrm{Sym}(\mathbb{N})$ is Borel, its isomorphism relation is Borel reducible to GI. The question is whether conversely, GI reduces to the isomorphism relation on the class. Mekler Mekler1985 (as noted in [FriedmanStanley1989, , 2.3]) established this for the countable groups that are nilpotent of class $2$, and of exponent a fixed odd prime. His method of coding graphs into such groups will be instrumental for our main result, Theorem 1.2. Paolini and Shelah PS showed it for countable torsion-free abelian groups. Most recently, Gao and Li GL did so for countable omnigenous locally finite groups (an uncountable class of groups, each generalising Hall’s universal locally finite group).
Open questions. An analytic equivalence relation on a Borel space is called universal analytic if each analytic equivalence relation is Borel reducible to it.
Question 1.5.
@@
We note that our original construction for the second step111see Version 1 of the arxiv preprint, available at arxiv.org/2512.12256.v1 used as the group $G_{n}$ in the inverse system $(G_{n},p_{n})_{n\in\mathbb{N}}$ the free Coxeter group generated by $V_{n}$, namely, the group freely generated by involutions $g_{v}$ for $v\in V_{n}$. That approach does not yield nilpotent-2 groups. However, it is somewhat simpler to verify, because the only elements of order 2 in $G_{n}$ are the conjugates of the generators, so that the elements of $V_{n}$ up to conjugacy can be recovered from $G_{n}$.
It is known that the relation of uniform homeomorphism on Polish metric spaces is universal analytic FLR .
@@
Is the equivalence relation of uniform homeomorphism on Polish ultrametric spaces universal analytic?
By the second step above, an affirmative answer to the second question would also answer the first in the affirmative. The preceding two questions will be discussed again in see Section 2.4.
Question 1.7.
@@
Notice that it follows from PS that graph isomorphism is a lower bound for the complexity of this classification problem, so an affirmative answer would fully determine the complexity.
@@ 2. Borel reduction of $\ell_{\infty}$ to uniform homeomorphism
We use the main result of CG , where it was shown that GI is Borel bireducible with the homeomorphism relation between 0-dimensional compact metric spaces. Let $\Phi$ denote a Borel witness to this reduction. Then for any graphs $\Gamma,\Gamma^{\prime}$ with vertex set $\mathbb{N}$, $\Gamma$ and $\Gamma^{\prime}$ are isomorphic if and only if $\Phi(\Gamma)$ and $\Phi(\Gamma^{\prime})$ are homeomorphic. Note that a Polish space is 0-dimensional if and only if it admits a compatible ultrametric. So there exists a compatible ultrametric $d_{\Gamma}$ for $\Phi(\Gamma)$. By the standard tools from descriptive set theory, we may find a Borel map giving the correspondence $\Phi(\Gamma)\mapsto d_{\Gamma}$.
@@ 2.1. Some facts on uniform homeomorphisms
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$. Then $\Phi$ is uniformly continuous if and only if
$\forall n\in\mathbb{N}\;\exists m\in\mathbb{N}\;\forall x,y\in[T]\;\bigl(x\!\upharpoonright\!m=y\!\upharpoonright\!m\;\rightarrow\;\Phi(x)\!\upharpoonright\!n=\Phi(y)\!\upharpoonright\!n\bigr).$
@@
Let $T,S$ be pruned subtrees of $\mathbb{N}^{<\mathbb{N}}$ and let $\Phi\colon[T]\to[S]$ be a uniform homeomorphism (see 1.4). If both $[T],[S]$ are uncountable, then $\Omega_{\Phi}$ and $\Omega_{\Phi^{-1}}$ are unbounded monotone functions from $\mathbb{N}$ to $\mathbb{N}$.
@@ 2.2. Definition of the reduction
The trees $\mathsf{S}_{1},\mathsf{S}_{2},\dots,\mathsf{S}_{k},\dots$ will be abstractly and inductively defined. Their exact realizations as subtrees of $\mathbb{N}^{<\mathbb{N}}$ will not be important. To begin, let $\mathsf{S}_{1}$ be a single branch. Thus the Cantor–Bendixson rank of $\mathsf{S}_{1}]$ is $1$. Let $\mathsf{S}_{2}$ be a pruned tree so that the Cantor–Bendixson derivative of $\mathsf{S}_{2}]$ is $\mathsf{S}_{1}]$. So the Cantor–Bendixson rank of $\mathsf{S}_{2}]$ is $2$. In general, let $\mathsf{S}_{k+1}$ be a pruned tree whose Cantor–Bendixson derivative is $\mathsf{S}_{k}]$. Then the Cantor–Bendixson rank of $\mathsf{S}_{k+1}$ is $k+1$. This finishes the definitions of $\mathsf{S}_{k}$ for $1\leqslant k<\mathbb{N}$.
@@
for any clopen set $V$ with $z\in V$, there exists a clopen set $U\subseteq V$ with $z\in U$ such that the $k$-th Cantor–Bendixson derivative of $U$ is the perfect kernel of $U$, and for any $i<k$, the $i$-th Cantor–Bendixson derivative of $U$ contains a perfect subset but is not itself perfect.
@@
Finally we are ready to define $\mathsf{T}_{x}$ given an $x\in\mathbb{N}^{\mathbb{N}}$. $\mathsf{T}_{x}$ is obtained from $\mathsf{S}_{\omega}$ by extending the node $\mathsf{t}_{k,2m+1}$ by a copy of $\mathsf{R}_{x(k)}$ for all $k,m\in\mathbb{N}$. This finishes the definition of $\mathsf{T}_{x}$. Figure 1 illustrates the definition of $\mathsf{T}_{x}$.
[…]
Figure 1. The construction of $\mathsf{T}_{x}$.